a wall has an area of 80 squared feet. what is the height of the wall?

Answer by checkley77(12844)

Many possible answers.
WIDTH*HEIGHT=80
8*10=80
10*8=80
5*16=80
16*5=80
7.5*10 2/3=80
10 2/3*7.5=80
6 2/3*12=80
12*6 2/3=80
4*20=80
20*4=80